# Multi-Preference Lambda-weighted Listwise DPO for Dynamic Preference Alignment

### v1

@@ Appendix B B Implementation of _compute_lambda_dpo_loss
B.1 Input Tensors and Log‑Probabilities
@@
    Load batch tensors:
@@
    Compute token-level and sequence-level log-probabilities:
    $\log\pi_{\theta}(y|x)$ $=\log\text{softmax}(\text{model}(x)[:,:-1])$
    mask $=\mathbb{1}_{\{y_{t}\neq\text{PAD}\}}$
    $\log\pi_{\theta}(y|x)$ $=\sum_{t}\texttt{mask}_{t}\cdot\log\pi_{\theta}(y_{t}|x_{t})$
    logits = model(input_ids=input_ids, attention_mask=attention_mask).logits
@@
    labels_shifted = labels[:, 1:].clone()
@@
    seq_log_probs = (label_log_probs * mask).sum(dim=1) / mask.sum(dim=1)
@@ B.2 Reference Model and Relative Reward
    Compute reference log-probabilities if available:
    $\log\pi_{\text{ref}}(y|x)=\begin{cases}\log\text{softmax}(\text{ref\_model}(x)[:,:-1]),&\text{if reference model exists}\\ 0,&\text{otherwise}\end{cases}$
    with torch.no_grad():
    ref_logits = self.ref_model(input_ids=input_ids, attention_mask=attention_mask).logits
    ref_log_probs = F.log_softmax(ref_logits[:, :-1], dim=-1)
    ref_seq_log_probs = (ref_label_log_probs * mask).sum(dim=1) / mask.sum(dim=1)
    else:
    ref_seq_log_probs = torch.zeros_like(seq_log_probs)
  2. 4.
    Compute relative reward and reshape:
    $r_{\theta}(x,y)$ $=\beta(\log\pi_{\theta}(y|x)-\log\pi_{\text{ref}}(y|x))$
@@ B.3 Listwise Loss by Preference Dimension
    For each dimension $i=1,2,3,4$:
    $\ell^{(i)}$ $=-\sum_{j}p^{*(i)}_{j}\cdot\log\text{softmax}(r^{(i)})_{j}$
@@ B.4 Lambda Weights and Final Loss
    Sample weights $\boldsymbol{\lambda}\sim\Delta^{4}$, and compute:

### v2

@@ Introduction
Conclusion
We presented Multi-Preference Lambda-weighted Listwise DPO, a generalization of Direct Preference Optimization (DPO) that enables alignment of language models with diverse and dynamic human preferences. By extending DPO from a pairwise to a listwise formulation and incorporating multiple preference dimensions via a lambda-weighted aggregation, our approach captures richer supervision signals and allows for controllable alignment without retraining.
Unlike traditional DPO, which assumes a single binary comparison per training instance, Lambda-DPO models the full distribution over ranked outputs, supporting scenarios where user intent is nuanced, multi-objective, or even conflicting. This enables practical applications such as personalized assistants that can adapt to individual user trade-offs (e.g., prioritizing helpfulness over conciseness), regulatory-sensitive deployments that balance safety and informativeness, and modular alignment pipelines where preferences may evolve over time.
Theoretically, our method is grounded in a KL-regularized RL objective, and empirically, we demonstrate competitive or superior performance to standard DPO across alignment benchmarks. By reducing estimation variance and providing interpretable preference control via the lambda vector, Lambda-DPO offers a scalable and expressive solution for real-world preference alignment.
In future work, we plan to explore more structured lambda scheduling, integration with curriculum learning, and extensions to conversational or multimodal alignment tasks.
@@ Appendix B B Implementation of _compute_lambda_dpo_loss
B.1 Input Tensors and Chunked Log-Probabilities
@@
    Load the batch tensors:
@@
    Set the chunk size and compute token/sequence log-probabilities:
    To reduce memory usage, we split the batch into chunks of size $C$, denoted by:
[…]
    chunk_size = self.lambda_dpo_chunk_size or input_ids.size(0)
    seq_log_probs_list = []
    ref_seq_log_probs_list = []
    for start in range(0, input_ids.size(0), chunk_size):
    end = start + chunk_size
    chunk_ids = input_ids[start:end]
    chunk_mask = attention_mask[start:end]
    chunk_labels = labels[start:end]
    logits = model(input_ids=chunk_ids, attention_mask=chunk_mask).logits
@@
    labels_shifted = chunk_labels[:, 1:].clone()
@@
    seq_lp = (label_log_probs * mask).sum(dim=1) / mask.sum(dim=1)
    seq_log_probs_list.append(seq_lp)
  3. 3.
    Compute reference log-probabilities (if applicable):
    $\log\pi_{\text{ref}}(y|x)=\begin{cases}\log\text{softmax}(\text{ref\_model}(x)),&\text{if ref model is used}\\ 0,&\text{otherwise}\end{cases}$
    with torch.no_grad():
    ref_logits = self.ref_model(input_ids=chunk_ids, attention_mask=chunk_mask).logits
    ref_log_probs = F.log_softmax(ref_logits[:, :-1], dim=-1)
    ref_lp = (ref_label_log_probs * mask).sum(dim=1) / mask.sum(dim=1)
    else:
    ref_lp = torch.zeros_like(seq_lp)
    ref_seq_log_probs_list.append(ref_lp)
    seq_log_probs = torch.cat(seq_log_probs_list, dim=0)
    ref_seq_log_probs = torch.cat(ref_seq_log_probs_list, dim=0)
@@ B.2 Reference Model and Relative Reward
    Compute the relative reward:
    $r_{\theta}(x,y)=\beta\left(\log\pi_{\theta}(y|x)-\log\pi_{\text{ref}}(y|x)\right)$
    Then reshape:
@@ B.3 Listwise Loss by Preference Dimension
    For each preference axis $i\in\{1,2,3,4\}$, compute the listwise loss:
    $\ell^{(i)}=-\sum_{j}p^{*(i)}_{j}\cdot\log\left(\frac{\exp(r^{(i)}_{j})}{\sum_{j^{\prime}}\exp(r^{(i)}_{j^{\prime}})}\right)$
@@ B.4 Lambda Weights and Final Loss
    Sample a random weight vector $\boldsymbol{\lambda}\sim\Delta^{4}$, and compute:
@@
Appendix C: Humorous Overview of the Added Classes
C.1 UltrafeedbackDatasetConverter
Input
A single UltraFeedback record. It contains:
  • a user prompt (instruction)
  • multiple completions, each with four ratings: helpfulness, honesty, instruction_following, and truthfulness.
Output
A dictionary with the prompt and, for every dimension, the raw scores plus a preference vector.
Example
Input:
instruction = "Explain photosynthesis."
completions = [
{"response": "...", "annotations": {"helpfulness": {"Rating": 4}}},
{"response": "...", "annotations": {"helpfulness": {"Rating": 2}}},
{"response": "...", "annotations": {"helpfulness": {"Rating": 5}}},
{"response": "...", "annotations": {"helpfulness": {"Rating": 3}}}
Output (simplified):
"_prompt": [{"role": "user", "content": "Explain photosynthesis."}],
"helpfulness": [
{"model": "?", "content": "...", "score": 4},
"_pi_target": {"helpfulness": [0.24, 0.09, 0.45, 0.22]}
C.2 ListwiseDatasetProcessor
Input
A dictionary produced by the converter that contains _prompt, the lists of responses for each dimension, and _pi_target.
Output
A flattened structure ready for training: for every sample there are 16 sequences of token IDs (one per response per dimension), matching labels and attention masks, and the list of probabilities.
Example
Input:
"_prompt": ...,
"helpfulness": [
{"content": "A"}, {"content": "B"}, {"content": "C"}, {"content": "D"}
"_pi_target": {"helpfulness": [0.24, 0.09, 0.45, 0.22], ...}
Output (illustrative shapes):
"input_ids": [seq1, seq2, ..., seq16], # each seq corresponds to a prompt+response
"labels": [...], # same layout as input_ids
"attention_mask": [...], # masks per sequence
"pi_target": [0.24, 0.09, 0.45, 0.22, ...] # flattened list of probabilities
C.3 ListwiseDataCollatorWithPadding
Input
A list of processed samples where each sample holds 16 sequences of token IDs and the flattened probability list.
Output
A padded batch containing tensors input_ids, attention_mask, labels, and an accompanying tensor of preference values aligned with the responses.
Example
Input:
# List of samples where each sample holds 16 sequences and
# a corresponding 16-element probability list
[example1, example2, ..., example_B] # B = batch size
Output:
"input_ids": Tensor of shape [B * 16, L],
"attention_mask": Tensor of shape [B * 16, L],
"labels": Tensor of shape [B * 16, L],
"pi_target": Tensor of shape [B * 16]
These three classes transform UltraFeedback data into tensors suitable for Lambda-weighted DPO training, while preserving the four-response structure and converting ratings into softmax preference vectors.